OpenStudy (anonymous):

Find the real or imaginary solutions of each equation by factoring. Did I do this right? And what's the difference between real and imaginary solutions? How do you tell the difference? x^3 – 64 = 0 x(x^2-64)=0 x=0 (x+8)(x-8)=0 x+8=0 x-8=0 x=0,-8,8

hartnn (hartnn):

64 is the cube of which number ?

hartnn (hartnn):

if it was x^3 -64x , then you would have been correct!

OpenStudy (anonymous):

Darn no it's not 64x What do I do to figure this out?

hartnn (hartnn):

use this formula \(a^3-b^3 = (a-b)(a^2+ab+b^2)\)

hartnn (hartnn):

here a = x b =4

OpenStudy (anonymous):

So x^3-64^3=(x-64)(x^2+64x+64^2)?

hartnn (hartnn):

the question is not 64^3 its 64, right ? 64 is the cube of what number?

OpenStudy (anonymous):

4

hartnn (hartnn):

yes so its actually x^3-4^3

hartnn (hartnn):

now use the formula

OpenStudy (anonymous):

x^3-4^3=(x-4)(x^2+4x+4^2)

hartnn (hartnn):

yes that = 0 so x-4 = 0 , x= .... (x^2+4x+4^2) = 0 can you solve this quadratic ?

OpenStudy (anonymous):

\[x=\frac{ \pm b \sqrt{b^2 -4ac} }{ 2 }\] This formula?

hartnn (hartnn):

yes \(\huge{x_{1,2}=\frac{-b \pm \sqrt{b^2-4ac}}{2a}}\)

hartnn (hartnn):

can you find a,b,c ?

OpenStudy (anonymous):

Also x=4 for x-4=0 Yeah I thought I clicked the a after the 2 lol.

hartnn (hartnn):

yes, x= 4 is correct :)

OpenStudy (anonymous):

a=1 b=4 c=16

hartnn (hartnn):

food! now find b^2-4ac = ... ?

hartnn (hartnn):

to answer your questions : And what's the difference between real and imaginary solutions? A : imaginary solutions have 'i' , imaginary unit \(i = \sqrt {-1}\) in them. they occur when we get a negative value inside square root. How do you tell the difference? A: if b^2 -4ac = positive, solutions are real if b^2-4ac is negative, then solutions are imaginary! so, here, whats b^2-4ac ?

OpenStudy (anonymous):

16-64=-48

hartnn (hartnn):

yes! so it came out to be negative! so your solutions will be complex :)

hartnn (hartnn):

now use the formula \(\huge{x_{1,2}=\frac{-b \pm \sqrt{b^2-4ac}}{2a}}\)

OpenStudy (anonymous):

\[x=\frac{ 4 \sqrt{-48} }{ 2 }\] \[x=\frac{ -4 \sqrt {-48} }{ 2 }\]

hartnn (hartnn):

you forgot \(\pm\) sign in between

OpenStudy (anonymous):

Indeed I did oops.

hartnn (hartnn):

so, could you simplify that ?

OpenStudy (anonymous):

Would it be x=-2 +,- sqrt -48 ?

hartnn (hartnn):

sqrt (-48) itself could be simplified... can you factor 48 ?

OpenStudy (anonymous):

No.

hartnn (hartnn):

\(48 = 16\times 3 \\ \sqrt {-48} = \sqrt {16} \sqrt {-3} = 4 \sqrt {-3}\) does that make sense ?

OpenStudy (anonymous):

Ahh okay! I know how to do that waqsn't sure that's what you meant.

hartnn (hartnn):

okk.. and \(4 \sqrt {-3} = 4 \sqrt {-1}\sqrt 3 = 4i \sqrt 3\)

OpenStudy (anonymous):

Okay confused. I don't know how to write this...

OpenStudy (anonymous):

I have to show my work and I'm stuck.

hartnn (hartnn):

i just used the fact that \(\sqrt{-1} = i\) nothing else...

OpenStudy (anonymous):

Yeah I'm stuck at the factoring part after I have \[x=\frac{ -4 \pm \sqrt {-48} }{ 2 }\] What was my next step after that?

hartnn (hartnn):

i already shown \(\sqrt {-48} = 4i \sqrt 3\)

OpenStudy (anonymous):

Yes I got that... So what \[x=-2 \pm 4i \sqrt{3}\]

hartnn (hartnn):